Please answer both questions and I will rate if both answers are correct.

Compute the following probabilities for the standard normal distribution Z B. P(-1.4 〈 Z 〈 0.15)- C. P(Z 〉-1.25)-
Answer #1

Answer - 1: From the z table output, the required areas will be:

A) Area = 0.4893

b) Area = 0.4789

c) Area = 0.8944

Answer - 2: A) P(X < 70000)

= P(z < (70000 - 76600)/3000)

= P(z < -2.2)

= 0.0139

B) P(X > 79600)

= P(z > (79600 - 76600)/3000)

= P(z > 1)

= 0.1587